# AdaAugment: A Tuning-Free and Adaptive Approach to Enhance Data Augmentation

### v2

@@ 3 AdaAugment
The policy determines the augmentation magnitudes $\boldsymbol{m}$ for the augmented data. Despite the randomness in the composition of each mini-batch during training, the magnitudes $\boldsymbol{m}$ operate on a per-sample basis, corresponding to each training sample. For simplicity, we denote the magnitudes of the current mini-batch data as $\boldsymbol{m}$, where the dimension of $\boldsymbol{m}$ equals the batch size, and each $m\in\boldsymbol{m}$ is strictly bounded within the interval $[0,1]$. When $m=0$, no augmentation is applied, while $m=1$ indicates the maximum magnitudes for the corresponding augmentation operations. In this way, magnitudes closer to 0 yield samples more similar to the original, whereas magnitudes closer to 1 produce more diverse data. The adaptively augmented sample $\tilde{\boldsymbol{x}}$ can be obtained by:
@@
where $e$ is a random augmentation operation and $m$ is determined based on the action policy. Therefore, Eq. (1) enables us to optimize the similarity-diversity preference of the augmented data. More importantly, this adjustment on the similarity-diversity preference reflects AdaAugment’s effectiveness in alleviating overfitting and underfitting risks.
@@
Consider a target classification model $f_{\theta}$, parameterized by $\theta$, an input sample $\boldsymbol{x}\in\mathbb{R}^{n}$, and $f_{\theta}(\boldsymbol{x})$ denotes the network output. Let $\mathcal{L}$ denote the cross-entropy loss, and $\mathcal{L}(f_{\theta}(\boldsymbol{x}),\boldsymbol{y})$ denote the loss item for the original sample $(\boldsymbol{x},\boldsymbol{y})$. Our proposed method aims to mitigate overfitting and underfitting risks by controlling the magnitudes of data augmentation operations based on feedback from the target model. To this end, we define three loss items based on our augmentation strategy. Firstly, $\mathcal{L}_{full}(f_{\theta}(\boldsymbol{x}^{+}),\boldsymbol{y})$ represents the loss of sample $\boldsymbol{x}$ with the maximal magnitude, i.e., $\boldsymbol{x}^{+}=e(m=1,\boldsymbol{x})$. Secondly, $\mathcal{L}_{none}(f_{\theta}(\boldsymbol{x}^{-}),\boldsymbol{y})$ denotes the loss for the non-augmented sample, i.e., $\boldsymbol{x}^{-}=e(m=0,\boldsymbol{x})=\boldsymbol{x}$. Lastly, based on Eq. (1), the loss for the adaptively augmented data is denoted as $\mathcal{L}_{ada}(f_{\theta}(\tilde{\boldsymbol{x}},\boldsymbol{y}))$, with the augmentation magnitude determined by the actor network. Inspired by the curriculum learning (Soviany et al., 2022), we formulate the reward function as follows:
@@
where $\lambda\in[0,1]$ is an adjusting factor, which is initialized at 1 and gradually decreases to 0 during the training process.
Policy Learning
The policy aims to determine the instance-level magnitudes for augmentation operations. For policy learning, we utilize the widely-used A2C algorithm (Mnih et al., 2016), which consists of an actor network $\theta_{a}$ and a critic network $\theta_{c}$, as depicted in Fig. 2. The actor network learns the policy, i.e., the probability distribution over actions given a particular state, $\pi(a\mid\boldsymbol{s})$. Meanwhile, the critic network aims to estimate the value associated with a particular state, denoted as $V^{\pi}(\boldsymbol{s})$.
Meanwhile, the loss function for updating $\theta_{c}$ is defined as:
@@
0:  dataset $\mathcal{D}$, batch size $B$, an augmentation space $\mathcal{E}$, cross-entropy loss $\mathcal{L}$, total number of epochs $T$
@@
In the realm of data augmentation, the augmentation magnitude is approximately proportional to the loss values of the augmented samples, i.e., $m\propto\mathcal{L}$. Therefore, the following inequality approximately holds: $\mathcal{L}_{none}\leq\mathcal{L}_{ada}\leq\mathcal{L}_{full}$, which suggests that $\mathcal{L}_{none}$ and $\mathcal{L}_{full}$ can serve as indicators of potential overfitting and underfitting risks associated with augmented samples, respectively. Meanwhile, these loss values continuously evolve as the training progresses.
@@
Regarding the reward function in Eq. (2), the term $\mathcal{L}_{full}-\mathcal{L}_{ada}$ holds significant importance during the initial training phases. Consequently, the policy network tends to employ relatively modest augmentation magnitudes, aiming to maximize the disparity between $\mathcal{L}_{full}$ and $\mathcal{L}_{ada}$. Therefore, more similar augmented samples can encourage models to capture broader features or patterns, accelerating rapid convergence during early training stages (Liu & Mirzasoleiman, 2022; Hou et al., 2023). As training advances, the focus shifts towards $\mathcal{L}_{ada}-\mathcal{L}_{none}$, prompting the policy network to apply greater augmentation magnitudes, trying to increase the difference between $\mathcal{L}_{ada}$ and $\mathcal{L}_{none}$. This yields more diverse augmented samples, mitigating overfitting risks. Thus, the methodology aligns with the curriculum learning principle (Soviany et al., 2022; Azad et al., 2023), wherein models initially learn patterns from similar samples and benefit from diverse samples in later training stages.
@@
Additionally, to investigate the similarity-diversity trade-off of AdaAugment throughout the training process, as depicted in Fig. 3, we evaluate the similarity and diversity measures (Yang et al., 2024) at various training stages. Specifically, we calculate the average magnitude values for all samples at the intervals of 50 epochs. Notably, it can be observed that the diversity measure increases and the similarity measure decreases gradually as the training progresses, which aligns with our theoretical analysis. This highlights AdaAugment’s inherent adaptability in adjusting DA strengths to harmonize with the evolving training dynamics of the target models.
Figure 3: Evolving magnitudes throughout the training process.
@@
Table 1: Test accuracy (%) on CIFAR-10/100 (average $\pm$ std). * means results reported in previous papers.
Dataset Method ResNet-18 ResNet-50 WRN-28-10 ShakeShake
CIFAR-10 baseline 95.28$\pm$0.14* 95.66$\pm$0.08* 95.52$\pm$0.11* 94.90$\pm$0.07*
CutMix 96.64$\pm$0.62* 96.81$\pm$0.10* 96.93$\pm$0.10* 96.47$\pm$0.07
Cutout 96.01$\pm$0.18* 95.81$\pm$0.17 96.92$\pm$0.09 96.96$\pm$0.09*
HaS 96.10$\pm$0.14* 95.60$\pm$0.15 96.94$\pm$0.08 96.89$\pm$0.10*
GridMask 96.38$\pm$0.17 96.15$\pm$0.19 97.23$\pm$0.09 96.91$\pm$0.12
RandomErasing 95.69$\pm$0.10 95.82$\pm$0.17 96.92$\pm$0.09 96.46$\pm$0.13*
AdvMask 96.44$\pm$0.15* 96.69$\pm$0.10* 97.02$\pm$0.05* 97.03$\pm$0.12*
AutoAugment 96.51$\pm$0.10* 96.59$\pm$0.04* 96.99$\pm$0.06 97.30$\pm$0.11
FAA 95.99$\pm$0.13 96.69$\pm$0.16 97.30$\pm$0.24 96.42$\pm$0.12
RandAugment 96.47$\pm$0.32 96.25$\pm$0.06 96.94$\pm$0.13* 97.05$\pm$0.15
DADA 95.58$\pm$0.06 95.61$\pm$0.14 97.30$\pm$0.13* 97.30$\pm$0.14*
TrivialAugment 96.28$\pm$0.10 97.07$\pm$0.08 97.18$\pm$0.11 97.30$\pm$0.10
Ours 96.75$\pm$0.06 97.34$\pm$0.13 97.66$\pm$0.07 97.41$\pm$0.06
CIFAR-100 baseline 77.54$\pm$0.19* 77.41$\pm$0.27* 78.96$\pm$0.25* 76.65$\pm$0.14*
CutMix 79.45$\pm$0.17 81.24$\pm$0.14 82.67$\pm$0.22 79.57$\pm$0.10
Cutout 78.04$\pm$0.10* 78.62$\pm$0.25 79.84$\pm$0.14 77.37$\pm$0.28
HaS 78.19$\pm$0.23 78.76$\pm$0.24 80.22$\pm$0.16 76.89$\pm$0.33
GridMask 75.23$\pm$0.21 78.38$\pm$0.22 80.40$\pm$0.20 77.28$\pm$0.38
AdvMask 78.43$\pm$0.18* 78.99$\pm$0.31* 80.70$\pm$0.25* 79.96$\pm$0.27*
RandomErasing 75.97$\pm$0.11* 77.79$\pm$0.32 80.57$\pm$0.15 77.30$\pm$0.18
AutoAugment 79.38$\pm$0.20 81.34$\pm$0.29 82.21$\pm$0.17 82.19$\pm$0.19
FAA 79.11$\pm$0.09 79.08$\pm$0.12 79.95$\pm$0.12 81.39$\pm$0.16
RandAugment 78.30$\pm$0.15 80.95$\pm$0.22 82.90$\pm$0.29* 80.00$\pm$0.29
DADA 78.28$\pm$0.22 80.25$\pm$0.28 82.50$\pm$0.26* 80.98$\pm$0.15
TrivialAugment 78.67$\pm$0.19 81.34$\pm$0.18 82.75$\pm$0.26 82.14$\pm$0.16
Ours 79.84$\pm$0.27 81.46$\pm$0.12 83.23$\pm$0.23 82.82$\pm$0.25
4 Experiments
Extensive experiments are conducted to evaluate the performance of our proposed method.
Experimental Settings
@@ 4 Experiments
To the best of our knowledge, image classification is the most widely used task for evaluating the effectiveness of DA methods (Lin et al., 2023a). We evaluate the effectiveness of our method across several benchmark datasets, including CIFAR10/100 (Krizhevsky & Hinton, 2010) and Tiny-ImageNet (Chrabaszcz et al., 2017). Additionally, we employ transfer learning to evaluate the efficacy of our method in enhancing model generalization performance. Moreover, we conduct effect visualization on MNIST (LeCun et al., 1998) and perform a convergence analysis of AdaAugment throughout the training process. A complexity analysis is incorporated to verify the computational efficiency and resource utilization of AdaAugment. Meanwhile, the effectiveness of AdaAugment on long-tail biased ImageNet-LT and Places-LT (Liu et al., 2019) is provided in Appendix LABEL:appendix:oltr.
Comparison with State-of-the-arts
@@
We compare our method with 10 most representative and commonly-used data augmentation methods, including CutMix (Yun et al., 2019), Cutout (DeVries & Taylor, 2017), HaS (Singh & Lee, 2017), GridMask (Chen et al., 2020), AdvMask (Yang et al., 2023), RandomErasing (Zhong et al., 2020), AutoAugment (Cubuk et al., 2019), Fast-AutoAugment (Lim et al., 2019), RandAugment (Cubuk et al., 2020), and TrivialAugment (Müller & Hutter, 2021).
Implementation Details
@@
The discount factor $\gamma$ in Eq. (3) and Eq. (4) is set as 0.99 in our method. We closely follow previous works (Yang et al., 2023; Müller & Hutter, 2021) in our setup. Specifically, all images are preprocessed by dividing each pixel value by 255 and normalized by the dataset statistics. For CIFAR-10 and CIFAR-100, we use ResNet architectures, including ResNet-18/50 (He et al., 2016), Wide-ResNet-28-10 (WRN-28-10) (Zagoruyko & Komodakis, 2016), and Shake-Shake-26x32 (Gastaldi, 2017). We train 1800 epochs with cosine learning rate decay for Shake-Shake using SGD with Nesterov Momentum and a learning rate of 0.01, a batch size of 256, $1e^{-3}$ weight decay. We train 300 epochs for all other networks using SGD with Nesterov Momentum and a learning rate of 0.1, a batch size of 128, a $5e^{-4}$ weight decay, and cosine learning rate decay. For Tiny-ImageNet, we resize the images into 64$\times$64, initialize models with ImageNet pre-trained weights, and then fine-tune models using various augmentations. The baseline approach involves solely padding and horizontal flipping. For a fair comparison, all methods are implemented with the same training configurations, conducting experiments across three independent random trials.
4.1 Results on CIFAR-10 and CIFAR-100
In Table 1, we evaluate the efficacy of AdaAugment on CIFAR-10 and CIFAR-100, employing various widely-used deep networks, including ResNet-18/50, Wide-ResNet-28-10, and ShakeShake-26-32. It can be observed that AdaAugment consistently demonstrates superior performance in improving the accuracy of these networks when compared to existing SOTA DA methods. Notably, our method demonstrates significant improvements for both datasets. For example, AdaAugment achieves improvements on the baseline model’s performance of ResNet-18, ResNet-50, Wide-ResNet-28-10, and ShakeShake-26-32 by 1.47%, 1.66%, 2.14%, and 2.51%, respectively. This superior performance of AdaAugment can be attributed to its adaptive adjustment of augmentation magnitudes based on the real-time feedback received from the target network during training. This adjustment effectively mitigates the risks of underfitting and overfitting, resulting in improved generalization capabilities.
4.2 Results on Tiny-ImageNet
In this section, we evaluate the efficacy of AdaAugment on the Tiny-ImageNet dataset using ResNet-18/50, Wide-ResNet-50-2 (WRN-50-2), and ResNext50 (Xie et al., 2017), as summarized in Table 2. It can be observed that AdaAugment demonstrates a significant enhancement in classification accuracy across various architectures, outperforming other widely employed DA methods by a large margin. Particularly, AdaAugment elevates the accuracy by 9.87%, 5.5%, 1.52%, and 1.16% on ResNet-18, ResNet-50, WRN-50-2, and ResNext-50, respectively. This performance enhancement underscores the efficacy of our novel adaptive augmentation mechanism. Thus, we demonstrate the superior efficacy and potential of AdaAugment for enhancing model performance.
@@ 4.2 Results on Tiny-ImageNet
Table 2: Image classification accuracy (%) on Tiny-ImageNet dataset is summarized in this table (average $\pm$ std).
@@
CutMix 64.09$\pm$0.30 76.41$\pm$0.27 82.32$\pm$0.46 81.31$\pm$1.00
Cutout 68.67$\pm$1.06 77.45$\pm$0.42 82.27$\pm$1.55 81.16$\pm$0.78
HaS 63.51$\pm$0.58 75.32$\pm$0.59 81.77$\pm$1.16 80.52$\pm$1.88
GridMask 62.72$\pm$0.91 77.88$\pm$2.50 82.25$\pm$1.47 81.05$\pm$1.33
AdvMask 65.29$\pm$0.20 78.84$\pm$0.28 82.87$\pm$0.55 81.38$\pm$1.54
RandomErasing 64.00$\pm$0.37 75.33$\pm$1.58 81.89$\pm$1.40 81.52$\pm$1.68
AutoAugment 67.28$\pm$1.40 75.29$\pm$2.40 79.99$\pm$2.20 81.28$\pm$0.33
FAA 68.15$\pm$0.70 75.11$\pm$2.70 82.90$\pm$0.92 81.04$\pm$1.92
RandAugment 65.67$\pm$1.10 75.87$\pm$1.76 82.25$\pm$1.02 80.36$\pm$0.62
DADA 70.03$\pm$0.10 78.61$\pm$0.34 83.03$\pm$0.18 81.15$\pm$0.34
TrivialAugment 69.97$\pm$0.96 78.41$\pm$0.39 82.16$\pm$0.32 80.91$\pm$2.26
Ours 71.25$\pm$0.64 79.11$\pm$1.51 83.07$\pm$0.78 81.92$\pm$0.29
4.3 Results on ImageNet

### v3

@@ 3 Linear (256, 1)
The policy determines the augmentation magnitudes $\boldsymbol{m}$ for the augmented data. Despite the randomness in the composition of each mini-batch during training, the magnitudes $\boldsymbol{m}$ operate on a per-sample basis, corresponding to each training sample. For simplicity, we denote the magnitudes of the current mini-batch data as $\boldsymbol{m}$, where the dimension of $\boldsymbol{m}$ equals the batch size, and each $m\in\boldsymbol{m}$ is strictly bounded within the interval $[0,1]$. When $m=0$, no augmentation is applied, while $m=1$ indicates the maximum magnitudes for the corresponding augmentation operations. Thus, magnitudes closer to 0 yield samples that are more similar to the original data, while magnitudes closer to 1 produce more diverse data. The adaptively augmented sample $\tilde{\boldsymbol{x}}$ can be obtained by:
@@
where $e$ is a random augmentation operation and $m$ is determined based on the action policy. Eq. (1) enables us to optimize the similarity-diversity preference of the augmented data. Thus, the policy network does not know the augmentation types and directions but just optimizes the magnitudes. More importantly, this adjustment on the similarity-diversity preference reflects AdaAugment’s effectiveness in accelerating model convergence in the early training stages and alleviating overfitting risks at later training stages.
@@
Consider a target classification model $f_{\theta}$, parameterized by $\theta$, an input sample $\boldsymbol{x}\in\mathbb{R}^{n}$, and $f_{\theta}(\boldsymbol{x})$ denotes the network output. Let $\mathcal{L}$ denote the cross-entropy loss, and $\mathcal{L}(f_{\theta}(\boldsymbol{x}),\boldsymbol{y})$ represent the loss item for the original sample $(\boldsymbol{x},\boldsymbol{y})$. Our proposed method aims to promote model fitting in the early training stages and alleviate overfitting at later training stages by controlling the magnitudes of data augmentation operations based on feedback from the target model. To this end, we define three loss terms based on our augmentation strategy. Firstly, $\mathcal{L}_{full}(f_{\theta}(\boldsymbol{x}^{+}),\boldsymbol{y})$ represents the loss for the sample $\boldsymbol{x}$ with the maximal augmentation magnitude, i.e., $\boldsymbol{x}^{+}=e(m=1,\boldsymbol{x})$. Secondly, $\mathcal{L}_{none}(f_{\theta}(\boldsymbol{x}^{-}),\boldsymbol{y})$ denotes the loss for the non-augmented sample, i.e., $\boldsymbol{x}^{-}=e(m=0,\boldsymbol{x})=\boldsymbol{x}$. Lastly, based on Eq. (1), the loss for the adaptively augmented data is denoted as $\mathcal{L}_{ada}(f_{\theta}(\tilde{\boldsymbol{x}},\boldsymbol{y}))$, with the augmentation magnitude determined by the actor network. Inspired by the curriculum learning [29], we formulate the reward function as follows:
@@
where $\lambda\in[0,1]$ is an adjusting factor, which is initialized at 1 and gradually decreases to 0 during the training process. Through this adaptive mechanism, we encourage the RL module to apply weaker DA at the beginning of training to accelerate model learning and progressively increase the DA magnitude in later stages to enhance sample diversity and mitigate the risk of overfitting.
@@
0:  dataset $\mathcal{D}$, batch size $B$, an augmentation space $\mathcal{E}$, cross-entropy loss $\mathcal{L}$, total number of epochs $T$.
@@
Policy Learning
The policy aims to determine the instance-level magnitudes for augmentation operations. For policy learning, we employ the widely-used A2C algorithm [64] due to its lightweight architecture. The details of the A2C architecture are presented in Table II, which consists of an actor network $\theta_{a}$ and a critic network $\theta_{c}$, which only contains several linear layers. The actor network learns the policy, i.e., the probability distribution over actions given a particular state, $\pi(a\mid\boldsymbol{s})$. Meanwhile, the critic network estimates the value associated with a particular state, denoted as $V_{\theta_{c}}(s)$.
Meanwhile, the loss function for updating $\theta_{c}$ is defined as:
Figure 3: Evaluation of the adaptively augmented datasets throughout the training process with CIFAR-10 using ResNet-50.
@@
In the realm of data augmentation, the augmentation magnitude is approximately proportional to the loss values of the augmented samples, i.e., $m\propto\mathcal{L}$. Therefore, the following inequality approximately holds: $\mathcal{L}_{none}\leq\mathcal{L}_{ada}\leq\mathcal{L}_{full}$, which suggests that $\mathcal{L}_{none}$ and $\mathcal{L}_{full}$ can serve as real-time indicators of potential overfitting and underfitting status associated with augmented samples, respectively. Meanwhile, these loss values continuously evolve as training progresses.
@@
Regarding the reward function in Eq. (2), the term $\mathcal{L}_{full}-\mathcal{L}_{ada}$ holds significant importance during the initial training phases. Consequently, the policy network tends to employ relatively modest augmentation magnitudes, aiming to maximize the disparity between $\mathcal{L}_{full}$ and $\mathcal{L}_{ada}$. Therefore, more similar augmented samples are encouraged initially, enabling models to capture broader features or patterns and facilitating rapid convergence during early training stages [67, 50]. As training advances, the focus shifts towards $\mathcal{L}_{ada}-\mathcal{L}_{none}$, prompting the policy network to apply greater augmentation magnitudes, trying to increase the difference between $\mathcal{L}_{ada}$ and $\mathcal{L}_{none}$. This yields more diverse augmented samples, thereby mitigating overfitting risks. Thus, the methodology aligns with the curriculum learning principle [29, 68], wherein models initially learn patterns from similar samples and subsequently benefit from diverse samples in later training stages.
@@
To further investigate the similarity-diversity trade-off of AdaAugment throughout the training process, we employ the similarity and diversity measures [9] for the augmented data at various training stages. Specifically, we calculate the average magnitude values for all samples at the intervals of 50 epochs. As depicted in Figure 3, it can be observed that the diversity measure increases and the similarity measure decreases gradually as training progresses, which aligns with our theoretical analysis. This highlights AdaAugment’s inherent adaptability in adjusting DA strengths to harmonize with the evolving training dynamics of the target models.
@@
TABLE III: Test accuracy (%) on CIFAR-10/100 (average $\pm$ std). * means results reported in previous papers.
Method CIFAR-10 CIFAR-100
ResNet-18 ResNet-50 WRN-28-10 ShakeShake ResNet-18 ResNet-50 WRN-28-10 ShakeShake
baseline 95.28$\pm$0.14* 95.66$\pm$0.08* 95.52$\pm$0.11* 94.90$\pm$0.07* 77.54$\pm$0.19* 77.41$\pm$0.27* 78.96$\pm$0.25* 76.65$\pm$0.14*
HaS [45] 96.10$\pm$0.14* 95.60$\pm$0.15 96.94$\pm$0.08 96.89$\pm$0.10* 78.19$\pm$0.23 78.76$\pm$0.24 80.22$\pm$0.16 76.89$\pm$0.33
FAA [46] 95.99$\pm$0.13 96.69$\pm$0.16 97.30$\pm$0.24 96.42$\pm$0.12 79.11$\pm$0.09 79.08$\pm$0.12 79.95$\pm$0.12 81.39$\pm$0.16
DADA [28] 95.58$\pm$0.06 95.61$\pm$0.14 97.30$\pm$0.13* 97.30$\pm$0.14* 78.28$\pm$0.22 80.25$\pm$0.28 82.50$\pm$0.26* 80.98$\pm$0.15
Cutout [14] 96.01$\pm$0.18* 95.81$\pm$0.17 96.92$\pm$0.09 96.96$\pm$0.09* 78.04$\pm$0.10* 78.62$\pm$0.25 79.84$\pm$0.14 77.37$\pm$0.28
CutMix [18] 96.64$\pm$0.62* 96.81$\pm$0.10* 96.93$\pm$0.10* 96.47$\pm$0.07 79.45$\pm$0.17 81.24$\pm$0.14 82.67$\pm$0.22 79.57$\pm$0.10
MADAug [50] 96.49$\pm$0.10 97.12$\pm$0.11 97.48$\pm$0.12 97.37$\pm$0.11 79.39$\pm$0.19 81.40$\pm$0.10 83.01$\pm$0.11 81.67$\pm$0.18
AdvMask [16] 96.44$\pm$0.15* 96.69$\pm$0.10* 97.02$\pm$0.05* 97.03$\pm$0.12* 78.43$\pm$0.18* 78.99$\pm$0.31* 80.70$\pm$0.25* 79.96$\pm$0.27*
GridMask [15] 96.38$\pm$0.17 96.15$\pm$0.19 97.23$\pm$0.09 96.91$\pm$0.12 75.23$\pm$0.21 78.38$\pm$0.22 80.40$\pm$0.20 77.28$\pm$0.38
AutoAugment [21] 96.51$\pm$0.10* 96.59$\pm$0.04* 96.99$\pm$0.06 97.30$\pm$0.11 79.38$\pm$0.20 81.34$\pm$0.29 82.21$\pm$0.17 82.19$\pm$0.19
RandAugment [22] 96.47$\pm$0.32 96.25$\pm$0.06 96.94$\pm$0.13* 97.05$\pm$0.15 78.30$\pm$0.15 80.95$\pm$0.22 82.90$\pm$0.29* 80.00$\pm$0.29
TeachAugment [25] 96.47$\pm$0.09 96.40$\pm$0.10 97.50$\pm$0.12 97.29$\pm$0.14 79.27$\pm$0.17 80.54$\pm$0.23 82.81$\pm$0.20 81.30$\pm$0.20
TrivialAugment [23] 96.28$\pm$0.10 97.07$\pm$0.08 97.18$\pm$0.11 97.30$\pm$0.10 78.67$\pm$0.19 81.34$\pm$0.18 82.75$\pm$0.26 82.14$\pm$0.16
RandomErasing [43] 95.69$\pm$0.10 95.82$\pm$0.17 96.92$\pm$0.09 96.46$\pm$0.13* 75.97$\pm$0.11* 77.79$\pm$0.32 80.57$\pm$0.15 77.30$\pm$0.18
AdaAugment 96.75$\pm$0.06 97.34$\pm$0.13 97.66$\pm$0.07 97.41$\pm$0.06 79.84$\pm$0.27 81.46$\pm$0.12 83.23$\pm$0.23 82.82$\pm$0.25
IV Experiments
@@
Experimental Settings Following previous work [16, 28, 23], we evaluate the effectiveness of our method across several benchmark datasets, including the coarse-grained classification datasets CIFAR-10/100 [69], Tiny-ImageNet [35], and ImageNet-1k [36]. We then employ transfer learning to evaluate the efficacy of our method in enhancing model generalization performance and transferability. Moreover, we demonstrate the efficacy of AdaAugment on long-tail classification datasets, including ImageNet-LT and Places-LT [37], as well as on several fine-grained datasets, including Oxford Flowers [39], Oxford-IIIT Pets [40], FGVC-Aircrafts [41], and Stanford Cars [38]. In addition, we conduct effect visualization and perform a convergence analysis of AdaAugment on CIFAR-10 throughout the training process. We also provide a complexity analysis of AdaAugment’s computational efficiency and resource utilization. Ablation studies are conducted to show how the components affect performance.
For baseline methods, AutoAugment and Fast-AutoAugment employ 16 operations, including Shear X/Y, Rotate, AutoContrast, Invert, Equalize, Solarize, Posterize, Contrast, Color, Brightness, Sharpness, Cutout, and Sample pairing. It randomly chooses from 25 augmentation policies on CIFAR-10/100 and 24 on ImageNet-1k. Each policy contains 2 augmentation operations. RandAugment contains 14 operations, which is the same as Table I, and it randomly selects $N$ operations from the augmentation space with fixed magnitudes of $M$. TeachAugment trains augmentation models to represent geometric augmentation and color augmentation, which can represent most operations employed in AutoAugment. TrivialAugment employs 21 augmentation operations. Besides the ones used in Table I, it also uses Cutout, Invert, Flip-X/Y, Sample pairing, Blur, and Smooth. Moreover, the probability of CutMix is 0.5 on CIFAR10/100 and 1.0 on ImageNet-1k. The probability of Cutout and AdvMask is 1.0.
@@
Comparison with State-of-the-arts We compare our method with 13 most representative and commonly-used data augmentation methods, including HaS [45], Fast-AutoAugment (FAA) [46], DADA [28], Cutout [14], CutMix [18], MADAug [50], AdvMask [16], GridMask [15], AutoAugment [21], RandAugment [22], TeachAugment [25], TrivialAugment [23], and RandomErasing [43].
@@
Implementation Details The discount factor $\gamma$ in Eq. (3) and Eq. (4) is set to 0.99 in our method, following [70, 71, 72]. We closely follow previous works [16, 23] in our setup. Specifically, all images are preprocessed by dividing each pixel value by 255 and normalized by the dataset statistics. For CIFAR-10 and CIFAR-100, we utilize ResNet-18/50 [73], Wide-ResNet-28-10 (WRN-28-10) [74], and Shake-Shake-26x32 (ShakeShake) [75]. We train 1800 epochs with cosine learning rate decay for Shake-Shake using SGD with Nesterov Momentum and a learning rate of 0.01, a batch size of 256, $1e^{-3}$ weight decay. We train 300 epochs for all other networks using SGD with Nesterov Momentum and a learning rate of 0.1, a batch size of 128, a $5e^{-4}$ weight decay, and cosine learning rate decay. For Tiny-ImageNet, we resize the images into 64$\times$64, initialize models with ImageNet pre-trained weights, and then fine-tune models using various augmentations. In each experiment, we perform three independent random trials. For ImageNet-1k, following [21, 23], ResNet-50 is used as the target network. We resize images into 224$\times$224, use the learning rate of 0.1, a batch size of 256, the Nesterov Momentum with a momentum parameter of 0.9, and a weight decay of 0.4. Note that because of the huge computational cost, the experiment in each case is performed once. The baseline approach involves solely padding and horizontal flipping. For a fair comparison, all methods are implemented with the same training configurations. If not specified, all experiments are conducted on an 8-NVIDIA-2080TI-GPU server.
@@
TABLE IV: Image classification accuracy (%) on Tiny-ImageNet dataset (average $\pm$ std).
@@
HaS [45] 63.51$\pm$0.58 75.32$\pm$0.59 81.77$\pm$1.16 80.52$\pm$1.88
FAA [46] 68.15$\pm$0.70 75.11$\pm$2.70 82.90$\pm$0.92 81.04$\pm$1.92
DADA [28] 70.03$\pm$0.10 78.61$\pm$0.34 83.03$\pm$0.18 81.15$\pm$0.34
Cutout [14] 68.67$\pm$1.06 77.45$\pm$0.42 82.27$\pm$1.55 81.16$\pm$0.78
CutMix [18] 64.09$\pm$0.30 76.41$\pm$0.27 82.32$\pm$0.46 81.31$\pm$1.00
MADAug [50] 70.16$\pm$0.76 78.62$\pm$0.32 82.38$\pm$0.42 81.41$\pm$1.26
AdvMask [16] 65.29$\pm$0.20 78.84$\pm$0.28 82.87$\pm$0.55 81.38$\pm$1.54
GridMask [15] 62.72$\pm$0.91 77.88$\pm$2.50 82.25$\pm$1.47 81.05$\pm$1.33
AutoAugment [21] 67.28$\pm$1.40 75.29$\pm$2.40 79.99$\pm$2.20 81.28$\pm$0.33
RandAugment [22] 65.67$\pm$1.10 75.87$\pm$1.76 82.25$\pm$1.02 80.36$\pm$0.62
TeachAugment [25] 70.05$\pm$0.57 70.56$\pm$0.44 82.95$\pm$0.13 81.39$\pm$0.97
TrivialAugment [23] 69.97$\pm$0.96 78.41$\pm$0.39 82.16$\pm$0.32 80.91$\pm$2.26
RandomErasing [43] 64.00$\pm$0.37 75.33$\pm$1.58 81.89$\pm$1.40 81.52$\pm$1.68
AdaAugment 71.25$\pm$0.64 79.11$\pm$1.51 83.07$\pm$0.78 81.92$\pm$0.29
IV-A Results on CIFAR-10 and CIFAR-100
In Table III, we evaluate the efficacy of AdaAugment on CIFAR-10 and CIFAR-100, employing various widely-used deep networks, such as ResNet-18/50 [73], Wide-ResNet-28-10 [74], and ShakeShake-26-32 [75]. It can be observed that AdaAugment consistently exhibits superior performance in improving the accuracy of these networks when compared to existing SOTA DA methods. Notably, our method demonstrates significant improvements for both datasets. For example, AdaAugment achieves improvements on the baseline model’s performance of ResNet-18, ResNet-50, Wide-ResNet-28-10, and ShakeShake-26-32 by 1.47%, 1.66%, 2.14%, and 2.51%, respectively. While on smaller models, the performance differences among different methods are marginal, using larger models can further highlight the effectiveness of AdaAugment. For instance, AdaAugment outperforms CutMix by nearly 1% on CIFAR-10 and 3.3% on CIFAR-100 with ShakeShake. This superior performance of AdaAugment can be attributed to its adaptive adjustment of augmentation magnitudes based on real-time feedback from the target network during training. This adjustment effectively mitigates the risks of overfitting, resulting in improved generalization capabilities. Additionally, larger models do not necessarily lead to better performance. The baseline performance of ResNet-50 underperforms that of ResNet-18 by applying simple augmentations. This is likely due to higher overfitting risks for the larger models. However, when more advanced augmentation methods are employed, ResNet-50 consistently outperforms ResNet-18, further validating the effectiveness of DA.
TABLE V: Image classification accuracy (%) on Tiny-ImageNet dataset with randomly initialized ResNet-50.
baseline HaS FAA DADA Cutout CutMix GridMask
49.73$\pm$.37 52.05$\pm$.12 52.83$\pm$.02 53.07$\pm$.28 50.89$\pm$.52 52.31$\pm$.20 50.37$\pm$.11
MADAug AA RA TeachAug TA RE AdaAugment
54.95$\pm$.24 53.37$\pm$.13 50.63$\pm$.53 53.97$\pm$.23 53.88$\pm$.03 50.74$\pm$.27 55.98$\pm$.12
IV-B Results on Tiny-ImageNet
In this section, we evaluate the efficacy of AdaAugment on the Tiny-ImageNet dataset using pre-trained ResNet-18/50, Wide-ResNet-50-2 (WRN-50-2), and ResNext50 [76], with the results summarized in Table IV. AdaAugment achieves significant improvements in classification accuracy across various architectures, outperforming other widely used DA methods by a substantial margin. Particularly, AdaAugment elevates the accuracy by 9.87%, 5.5%, 1.52%, and 1.16% on ResNet-18, ResNet-50, WRN-50-2, and ResNext-50, respectively. Furthermore, Table V presents the results of training ResNet-50 on Tiny-ImageNet from scratch. Even in this scenario, AdaAugment delivers superior performance, achieving an improvement of over 1%. These results highlight the effectiveness of our adaptive augmentation mechanism in enhancing both pre-trained and from-scratch training setups. Thus, we demonstrate the superior efficacy and potential of AdaAugment for enhancing model performance.